# On the analysis of a higher-order Lotka-Volterra model: an application of $\mathcal{S}$-tensors and the polynomial complementarity problem

### v1

@@ II-D Polynomial complementarity problem
Lemma 5 (Theorem 7.1 [18], Proposition 2.1 [19])
@@
(a) $\operatorname{SOL}\left(F^{\infty},0\right)=\{0\}$.
(b) For any bounded set $\Omega\subset\mathbb{R}^{n},\bigcup_{q\in\Omega}\operatorname{SOL}(F,q)$ is bounded.
Then, $(a)\Rightarrow(b)$. The reverse implication holds when $F$ is homogeneous.
@@
Lemma 6 (Theorem 7.2 [18], Theorem 5.1 [19])
@@
(a) $\operatorname{SOL}\left(F^{\infty},0\right)=\{0\}=\operatorname{SOL}\left(F^{\infty},d\right)$.
(b) $\operatorname{SOL}\left(F^{\infty},0\right)=\{0\}=\operatorname{SOL}(F,d)$.
@@
The polynomial complementarity problem is said to have the property of global uniqueness and solvability (denoted by GUS-property), if and only if it has a unique solution for every $q\in\mathbb{R}^{n}$. A mapping $F:\mathbb{S}\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is said to be a $P$-function on $\mathbb{S}$, if and only if $\max_{i=1,\cdots,n}\left(x_{i}-y_{i}\right)\left[F_{i}(x)-F_{i}(y)\right]>0$ for all $x,y\in\mathbb{S}$ with $x\neq y$. The following lemma describes when the PCP has the GUS-property.
Lemma 7 (Theorem 7.4 [18], Theorem 6.1 [19])
@@
Now, equipped with the knowledge introduced above, we consider a special case, where $F(x)=Ax+Bx^{2}$. The polynomial complementarity problem now becomes a quadratic complementarity problem as follows:
$x\geq 0,\quad Bx^{2}+Ax+q\geq 0,\quad\text{ and }\quad x^{\top}\left(Bx^{2}+Ax+q\right)=0.$ (13)
@@
$r_{i}({A})_{-}:=\sum_{a_{ii_{2}i_{3}\cdots i_{m}}<0,\left(i_{2},i_{3},\ldots,i_{m}\right)\neq(i,i,\ldots,i)}\left|a_{ii_{2}i_{3}\cdots i_{m}}\right|;$ (15)
@@
A tensor ${A}=\left(a_{i_{1}i_{2}\cdots i_{m}}\right)\in\mathbb{R}^{[m,n]}$ is said to be a generalized row strictly diagonally dominant if and only if for all $i\in[n]$, $\left|a_{ii\cdots i}\right|-r_{i}({A})_{-}>0,$ where $r_{i}({A})_{-}$is defined by (15). If ${A}$ additionally satisfies $a_{ii\cdots i}>0$ for all $i\in[n]$, then we call ${A}$ a generalized row strictly diagonally dominant tensor with all positive diagonal entries.
Remark 3
@@
It refers to the set of index where the $i$-th component of $q$ is strictly negative.
Lemma 8
@@
Then we have $x_{k}\left(Bx^{2}+Ax+q\right)_{k}>0,$ which contradicts the fact that $x$ solves (13). Thus, it holds that $k\in$ $\Omega(q)$. Moreover, we have $\left(Bx^{2}+Ax\right)_{k}+q_{k}=0$ because $x_{k}>0$ and $x_{k}\left(Bx^{2}+Ax+q\right)_{k}=0$. ∎
@@
Proof:
The proof is similar to [28]. According to Lemma 8, there exists $k\in\Omega(q)$ such that $x_{k}=\|x\|_{\infty}>0$ and $\left(Bx^{2}+Ax\right)_{k}+q_{k}=0$. Thus, we have
@@
III General Higher-order Lotka-Volterra model
A purely competitive higher-order Lotka-Volterra model is proposed in [11]. Here, we consider both, competition and cooperation, among species. Inspired by [11], we propose here a general higher-order Lotka-Volterra model:
@@ III General Higher-order Lotka-Volterra model
Proof:
It is sufficient to notice that if $x_{i}=0$, $\dot{x}_{i}=0$. Once the trajectory enters the boundary, since its derivative becomes zero, it will never leave the first orthant. ∎
@@
We have the following result as a direct consequence of Theorem 1.
Corollary 2
System (19) has one unique positive equilibrium if $-A,\,-B$ are both $\mathcal{S}$-tensors associated with the same positive vector $v$, i.e., $(-A)v>0$, $(-B)v^{2}>0$.
@@
Proof:
It is straightforward to check that the equilibria of the system (19) with the restriction $1+\sum_{j}a_{ij}x^{*}_{j}+\sum_{j,k}b_{ijk}x^{*}_{j}x^{*}_{k}\leq 0$ yields the solution set of $\operatorname{SOL}\left(F,-\mathbf{1}\right)$ with $F=-Ax-Bx^{2}$, while the system (19) with the restriction $1+\sum_{j}a_{ij}x^{*}_{j}+\sum_{j,k}b_{ijk}x^{*}_{j}x^{*}_{k}\geq 0$ yields the solution set of $\operatorname{SOL}\left(-F,\mathbf{1}\right)$. ∎
Remark 5
Notice that the solution of the polynomial complementarity problem satisfies $01+\sum_{j}a_{ij}x^{*}_{j}+\sum_{j,k}b_{ijk}x^{*}_{j}x^{*}_{k}<$ 0 if $x_{i}^{*}=0$, $01+\sum_{j}a_{ij}x^{*}_{j}+\sum_{j,k}b_{ijk}x^{*}_{j}x^{*}_{k}=$ 0 if $x_{i}^{*}>0$. In this paper, we will then show that (21) is related to the stability of some boundary equilibria (equilibria on the boundary of the first orthant). Thus, the polynomial complementarity problem will help us to find potential stable equilibria. Moreover, by further utilizing the result of Lemmas 5-7, we can know when the equilibrium set satisfying (21) is non-empty and bounded, or when it is a nonempty singleton. It is known that a classical Lotka-Volterra model is related to the linear complementarity problem [7, 17]. Here, we show the relationship between PCP and the higher-order Lotka-Volterra model.
@@
Proof:
It is straightforward to see that $\mathbf{0}$ is always a solution of $R\operatorname*{\mathsf{diag}}(x)\left(\mathbf{1}+Ax+Bx^{2}\right)$. For the equilibrium point $\mathbf{0}$, the corresponding Jacobian matrix is $\mathbf{J}_{(\mathbf{0})}=\operatorname*{\mathsf{diag}}((r_{1},\cdots,r_{n})^{\top})$. Since $r_{i}>0$, the equilibrium point $\mathbf{0}_{n}$ is unstable. ∎
@@
where $F(x)=(F_{1}(x),\cdots,F_{n}(x))^{\top}$ and $F_{i}(x)=r_{i}+r_{i}\sum^{n}_{j=1}a_{ij}x_{j}+r_{i}\sum^{n}_{j,k=1}b_{ijk}x_{j}x_{k}$. According to [31, Theorem 5], we have the following result as a corollary. The definition and the concept are consistent with the paper [31]. For a set $\mathcal{M}$ with boundary, we denote the boundary as $\partial\mathcal{M}$, and the interior $\operatorname{Int}(\mathcal{M})\triangleq\mathcal{M}\backslash\partial\mathcal{M}$. The concept of ”pointing inward” is defined the same as in [31, Definition 1].
Corollary 4
@@
As a corollary of [32, Theorem 4], we have the following:
Corollary 5
@@
We just need to show that the conditions in Corollary 5 are satisfied. More precisely, one needs to show the Jacobian is bounded by a matrix $G$ and all the leading principal minors of $-G$ are positive. We have that $\frac{\partial F_{i}(x)}{\partial x_{i}}=r_{i}a_{ii}+r_{i}\sum_{k}b_{ijk}x_{k}+r_{i}\sum_{k}b_{ikj}x_{k}\leq r_{i}a_{ii}=G_{ii}<0$. On the other hand, $\left|\frac{\partial F_{i}(x)}{\partial x_{j}}\right|\leq r_{i}\{\sum_{j\neq i}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(|b_{ijk}|+|b_{ikj}|)\}=G_{ij}.$ The condition $-d_{i}a_{ii}>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$ guarantees that all the leading principal minors of $-G$ are positive, according to the Lemma 6 of [31]. Then, we can utilize corollary 5. ∎
Remark 6
@@
For the condition $-d_{i}a_{ii}>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$, if all $d_{i}=1,\forall i$, the matrix $-A$ is diagonally dominant with all positive diagonal elements, which is an $\mathcal{S}-$tensor. If $B$ is nonpositive, then $-B$ is also an $\mathcal{S}-$tensor. Thus, Corollary 6 is compatible with Theorem 1. We can firstly use Theorem 1 to see whether there exists a positive equilibrium and then use corollary 6 to determine the uniqueness and global stability.
Next, we propose the most general result on the global stability by using the tensor properties.
Theorem 5
Consider the system (19) with $A,B$ both strictly diagonally dominant tensor with negative diagonal entries. The system has a unique positive equilibrium and it is globally asymptotically stable.
Proof:
The equation for a positive equilibrium is $-Ax-Bx^{2}=\mathbf{1}$. It is straightforward to know that $-A,-B$ are both $\mathcal{S}$-tensors with a positive vector $\mathbf{1}$, i.e. $-A\mathbf{1}>\mathbf{0},-B\mathbf{1}^{2}>\mathbf{0}$, according to the definition of a strictly diagonally dominant tensor. From Theorem 1, there is a unique positive equilibrium.
Then, define $V_{m}=\max_{i}x_{i}^{2}$ and $m=\arg\max_{i}x_{i}^{2}$.
$\begin{split}\dot{V}_{m}&=2x_{i}\dot{x}_{m}=2x_{i}\left((Ax+Bx^{2})_{m}+1\right)\\ &=2x_{i}(\sum^{n}_{j=1}a_{mj}x_{j}+\sum^{n}_{j,k=1}b_{mjk}x_{j}x_{k}+1)\\ &\leq 2x_{i}(\sum^{n}_{j=1}a_{mj}x_{m}+\sum^{n}_{j,k=1}b_{mjk}x_{m}^{2}+1).\end{split}$ (26)
We can observe that if $x_{m}=1$, $A\mathbf{1}<\mathbf{0},B\mathbf{1}^{2}<\mathbf{0}$. Furthermore, if $x_{m}=\delta>0$, $A(\delta\mathbf{1})=\delta(A\mathbf{1}),B(\delta\mathbf{1}^{2})=\delta^{2}(B\mathbf{1}^{2})$. If $x_{m}$ is sufficiently big, $\dot{V}_{m}$ is negative. Thus, $V_{m}$ is upper bounded. Since $V_{m}$ is also lower bounded by zero, $V_{m}$ must have a finite limit. Because of the boundedness, we just need to deal with a bounded interval. Since $0\leq x_{i}\leq x_{m}$ for any $i$, $x_{i}$ has also a finite limit. Moreover, $\dot{x}_{i}$ is a polynomial function and thus uniformly continuous on a bounded interval. From Barbalat’s lemma, all $\dot{x}_{i}$ converges to zero and this corresponds to either the origin or the positive equilibrium. Since the origin is always unstable, it is excluded. ∎
Remark 7
We clearly see that the result of Theorem 5 is more general than corollary 4 and corollary 6.
@@
Proof:
The unperturbed system can be represented as $\dot{x}=g(x)$ and the perturbed system as $\dot{x}=G(x,\epsilon)$, with $G(x,0)=g(x)$. Let $x^{*}$ be a hyperbolic equilibrium of the unperturbed system. By definition of the equilibrium point, $G(x^{*},0)=0$ and $\frac{\partial G}{\partial x}(x^{*},0)=\frac{\partial g}{\partial x}(x^{*})$. Due to the hyperbolicity of $x^{*}$, $\frac{\partial G}{\partial z}(x^{*},0)=\frac{\partial g}{\partial x}(x^{*})$ has a nonvanishing determinant. By the implicit function theorem, there is a unique equilibrium in the neighborhood of $x^{*}$ for sufficiently small $\epsilon$. This equilibrium is also hyperbolic because of the continuous dependence of the eigenvalues of $\frac{\partial G}{\partial x}$ on $\epsilon$. Thus, the local stability of the equilibrium persists. ∎
Next, we treat the pairwise interaction as a perturbation. Let us now consider $0<\epsilon\ll 1$ and the perturbed system:
$\dot{x}=R\operatorname*{\mathsf{diag}}(x)\left(\mathbf{1}+\epsilon Ax+Bx^{2}\right).$ (28)
The unperturbed system is now
$\dot{x}=R\operatorname*{\mathsf{diag}}(x)\left(\mathbf{1}+Bx^{2}\right).$ (29)
As a corollary of Lemma 2, we have the following result.
Corollary 7
Consider the unperturbed system (29). If $B$ is an $\mathcal{S}$-tensor satisfying $C(ii)$ in Lemma 2, then the system (29) has a unique positive equilibrium.
We can always check the local stability of the unique positive equilibrium and whether it is hyperbolic by calculating the eigenvalues of the Jacobian at the unique positive equilibrium. We now provide a similar result to Theorem 6.
Theorem 7
Consider the perturbed system (28). If the unperturbed system (29) has a hyperbolic equilibrium point $x^{*}$, then the perturbed system (28) also has a hyperbolic equilibrium point $\tilde{x}^{*}$ in the vicinity of the hyperbolic equilibrium point $x^{*}$. Furthermore, if $x^{*}$ is locally stable, then $\tilde{x}^{*}$ is locally stable. Otherwise, if $x^{*}$ is unstable, then $\tilde{x}^{*}$ is unstable.
In this section, we have provided abundant results regarding the existence of a positive equilibrium. Notice that there may also be some boundary equilibria (besides the origin). However, these non-zero boundary equilibria can be considered as a positive equilibrium of a subsystem, the one restricted to the boundary. Moreover, we can also use the solution of the polynomial complementarity problem to find some potential stable boundary equilibria. Thus, all the results and techniques in this section are also applicable to that case.
IV Cooperative Lotka-Volterra Model

### v2

@@ 2.4 Polynomial complementarity problem
Lemma 2.5 (Theorem 7.1 [18], Proposition 2.1 [19]).
@@
  • (a)
    $\operatorname{SOL}\left(F^{\infty},0\right)=\{0\}$.
  • (b)
    For any bounded set $\Omega\subset\mathbb{R}^{n},\bigcup_{q\in\Omega}\operatorname{SOL}(F,q)$ is bounded.
Then, (a) implies (b). The reverse implication holds when $F$ is homogeneous.
@@
Lemma 2.6 (Theorem 7.2 [18], Theorem 5.1 [19]).
@@
  • (a)
    $\operatorname{SOL}\left(F^{\infty},0\right)=\{0\}=\operatorname{SOL}\left(F^{\infty},d\right)$.
  • (b)
    $\operatorname{SOL}\left(F^{\infty},0\right)=\{0\}=\operatorname{SOL}(F,d)$.
@@
The polynomial complementarity problem is said to have the property of global uniqueness and solvability (denoted by GUS-property) if and only if it has a unique solution for every $q\in\mathbb{R}^{n}$. A mapping $F:\mathbb{S}\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is said to be a $P$-function on $\mathbb{S}$, if and only if $\max_{i=1,\cdots,n}\left(x_{i}-y_{i}\right)\left[F_{i}(x)-F_{i}(y)\right]>0$ for all $x,y\in\mathbb{S}$ with $x\neq y$. The following lemma describes when the PCP has the GUS-property.
Lemma 2.7 (Theorem 7.4 [18], Theorem 6.1 [19]).
@@
Now, equipped with the preliminaries introduced above, we consider the special case where $F(x)=Ax+Bx^{2}$. The polynomial complementarity problem now becomes a quadratic complementarity problem as follows:
$\begin{split}x\geq\mathbf{0}\\ Bx^{2}+Ax+q\geq\mathbf{0}\\ x^{\top}\left(Bx^{2}+Ax+q\right)=0.\end{split}$ (13)
@@
$r_{i}({A})_{-}:=\sum_{a_{ii_{2}i_{3}\cdots i_{m}}<0,\left(i_{2},i_{3},\ldots,i_{m}\right)\neq(i,i,\ldots,i)}\left|a_{ii_{2}i_{3}\cdots i_{m}}\right|.$ (15)
@@
The tensor ${A}=\left(a_{i_{1}i_{2}\cdots i_{m}}\right)\in\mathbb{R}^{[m,n]}$ is said to be a generalized row strictly diagonally dominant tensor if and only if for all $i\in[n]$, $\left|a_{ii\cdots i}\right|-r_{i}({A})_{-}>0,$ where $r_{i}({A})_{-}$is defined by (15). If ${A}$ additionally satisfies $a_{ii\cdots i}>0$ for all $i\in[n]$, then we call ${A}$ a generalized row strictly diagonally dominant tensor with all positive diagonal entries.
Remark 2.8.
@@
The set $\Omega(q)$ refers to the set of indices where the $i$-th component of $q$ is strictly negative.
Lemma 2.9.
@@
Then we have $x_{k}\left(Bx^{2}+Ax+q\right)_{k}>0,$ which contradicts the fact that $x$ solves (13). Thus, it holds that $k\in$ $\Omega(q)$. Moreover, we have $\left(Bx^{2}+Ax\right)_{k}+q_{k}=0$ because $x_{k}>0$ and $x_{k}\left(Bx^{2}+Ax+q\right)_{k}=0$.
@@
Proof 2.11.
The proof is similar to [30]. According to Lemma 2.9, there exists $k\in\Omega(q)$ such that $x_{k}=\|x\|_{\infty}>0$ and $\left(Bx^{2}+Ax\right)_{k}+q_{k}=0$. Thus, we have
@@
3 General Higher-order Lotka-Volterra model
A purely competitive higher-order Lotka-Volterra model is proposed in [11]. Here, we consider both, competition and cooperation, among species. Inspired by [11], we study here a general higher-order Lotka-Volterra model:
@@ 3 General Higher-order Lotka-Volterra model
Proof 3.3.
It is sufficient to notice that if $x_{i}=0$, $\dot{x}_{i}=0$. Once the trajectory enters the boundary, since its derivative becomes zero, it will never leave the first orthant.
@@
Proof 3.4.
The proof is straightforward and can be done by observation.
Remark 3.5.
Notice that the solution of the polynomial complementarity problem satisfies $01+\sum_{j}a_{ij}x^{*}_{j}+\sum_{j,k}b_{ijk}x^{*}_{j}x^{*}_{k}<$ 0 if $x_{i}^{*}=0$, and $01+\sum_{j}a_{ij}x^{*}_{j}+\sum_{j,k}b_{ijk}x^{*}_{j}x^{*}_{k}=$ 0 if $x_{i}^{*}>0$. In this paper, we will show that (21) is related to the stability of some boundary equilibria (equilibria on the boundary of the first orthant). Thus, the polynomial complementarity problem will help us to find potential stable equilibria. Moreover, by further utilizing the result of Lemmas 2.5-2.7, we can know when the equilibrium set satisfying (21) is non-empty and bounded, or when it is a nonempty singleton. It is known that a classical Lotka-Volterra model is related to the linear complementarity problem [7, 17]. Here, we show the relationship between PCP and the higher-order Lotka-Volterra model.
@@
Proof 3.6.
It is straightforward to see that $\mathbf{0}$ is always a solution of $R\operatorname*{\mathsf{diag}}(x)\left(\mathbf{1}+Ax+Bx^{2}\right)$. For the equilibrium point $\mathbf{0}$, the corresponding Jacobian matrix is $\mathbf{J}_{(\mathbf{0})}=\operatorname*{\mathsf{diag}}((r_{1},\cdots,r_{n})^{\top})$. Since $r_{i}>0$, the equilibrium point $\mathbf{0}_{n}$ is unstable.
@@
where $F(x)=(F_{1}(x),\cdots,F_{n}(x))^{\top}$ and $F_{i}(x)=r_{i}+r_{i}\sum^{n}_{j=1}a_{ij}x_{j}+r_{i}\sum^{n}_{j,k=1}b_{ijk}x_{j}x_{k}$. We have the following result as a corollary of [33, Theorem 5]. The definition and the concept are consistent with the paper [33]. For a set $\mathcal{M}$ with boundary, we denote the boundary as $\partial\mathcal{M}$, and the interior $\operatorname{Int}(\mathcal{M})\triangleq\mathcal{M}\backslash\partial\mathcal{M}$. The concept of “pointing inward” is defined as in [33, Definition 1].
Corollary 7.
@@
As a corollary of [34, Theorem 4], we have the following:
Corollary 8.
@@
We just need to show that the conditions in Corollary 8 are satisfied. More precisely, one needs to show the Jacobian is bounded by a matrix $G$ and all the leading principal minors of $-G$ are positive. We have that $\frac{\partial F_{i}(x)}{\partial x_{i}}=r_{i}a_{ii}+r_{i}\sum_{k}b_{iki}x_{k}+r_{i}\sum_{k}b_{iik}x_{k}\leq r_{i}a_{ii}=G_{ii}<0$. On the other hand, $\left|\frac{\partial F_{i}(x)}{\partial x_{j}}\right|\leq r_{i}\{\sum_{j\neq i}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(|b_{ijk}|+|b_{ikj}|)\}=G_{ij}.$ The condition $-d_{i}a_{ii}>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$ guarantees that all the leading principal minors of $-G$ are positive, according to the Lemma 6 of [33]. Then, we can utilize corollary 8.
Remark 3.8.
@@
For the condition $-d_{i}a_{ii}>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$, it means that the diagonal entries of pairwise interaction dominates the whole dynamics. If all $d_{i}=1,\forall i$, the matrix $-A$ is diagonally dominant with all positive diagonal elements, which is an $\mathcal{S}-$tensor. If $B$ is nonpositive, then $-B$ is also an $\mathcal{S}-$tensor. Thus, Corollary 9 is compatible with Theorem 1.
A slightly more general result is shown by the following.
Corollary 10.
Consider the system (19) with $b_{iii}<0$ and $a_{ii}<0$ for all $i$. Suppose that there exist constants $\hat{R},\,\epsilon>0$ such that $\mathcal{W}\triangleq\left\{x:\|x\|\leq\hat{R},x_{i}\geq\epsilon\quad\forall i=1,\ldots,n\right\}$ is a positive invariant set. If there exists some coefficients $d_{i}$, $i=1,\cdots,n$ such that $-d_{i}(a_{ii}+b_{iii}\epsilon)>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{(j,k)\neq(i,i)}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$ for all $i$, then the positive equilibrium $x^{*}$, if it exists, is unique and globally asymptotically stable.
Proof 3.9.
We just need to show that the conditions in Corollary 8 are satisfied. More precisely, one needs to show the Jacobian is bounded by a matrix $G$ and all the leading principal minors of $-G$ are positive. We have that $\frac{\partial F_{i}(x)}{\partial x_{i}}=r_{i}a_{ii}+r_{i}\sum_{k}b_{iki}x_{k}+r_{i}\sum_{k}b_{iik}x_{k}\leq r_{i}(a_{ii}+b_{iii}\epsilon+\sum_{b_{iki}>0}b_{iki}\hat{R}+\sum_{b_{iki}>0}b_{iik}\hat{R})=G_{ii}<0$. On the other hand, $\left|\frac{\partial F_{i}(x)}{\partial x_{j}}\right|\leq r_{i}\{\sum_{j\neq i}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(|b_{ijk}|+|b_{ikj}|)\}=G_{ij}.$ The condition $-d_{i}(a_{ii}+b_{iii}\epsilon)>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{(j,k)\neq(i,i)}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$ guarantees that all the leading principal minors of $-G$ are positive, according to the Lemma 6 of [33]. Then, we can utilize corollary 8.
We can firstly use Theorem 1 to see whether there exists a positive equilibrium and then use Corollary 9 and 10 to determine the uniqueness and global stability. That is to first use the following result.
Theorem 11.
System (19) has a unique equilibrium point if $-A$ and $-B$ are all $\mathcal{S}-$tensors.
Proof 3.10.
The equation for a positive equilibrium is $-Ax-Bx^{2}=\mathbf{1}$. From Theorem 1, there is a unique positive equilibrium.
Remark 3.11.
The result is obviously stronger than Corollary 7. Here, we do not need a condition of a bounded positively invariant set. Recall that a strictly diagonally dominant tensor with all positive diagonal entries is $\mathcal{S}-$tensors. One observation from Theorem 11 and Corollary 10 is that if the negative diagonal entries of $a_{ii},b_{iii}$ are sufficiently large (in absolute value) compared with the off-diagonals, then the system has a unique globally asymptotically stable equilibrium. Corollary 10 is conservative and diagonal entries of $a_{ii},b_{iii}$ maybe sometimes huge. In the latter result of the Theorem 15 and 24, we study a much more relaxed condition under some specific scenarios.
@@
Proof 3.12.
The unperturbed system can be represented as $\dot{x}=g(x)$ and the perturbed system as $\dot{x}=G(x,\epsilon)$, with $G(x,0)=g(x)$. Let $x^{*}$ be a hyperbolic equilibrium of the unperturbed system. By definition of the equilibrium point, $G(x^{*},0)=0$ and $\frac{\partial G}{\partial x}(x^{*},0)=\frac{\partial g}{\partial x}(x^{*})$. Due to the hyperbolicity of $x^{*}$, $\frac{\partial G}{\partial z}(x^{*},0)=\frac{\partial g}{\partial x}(x^{*})$ has a nonvanishing determinant. By the implicit function theorem, there is a unique equilibrium in the neighborhood of $x^{*}$ for sufficiently small $\epsilon$. This equilibrium is also hyperbolic because of the continuous dependence of the eigenvalues of $\frac{\partial G}{\partial x}$ on $\epsilon$. Thus, the local stability of the equilibrium persists.
In this section, we have provided abundant results regarding the existence of a positive equilibrium and its stability in some cases. Notice that there may also be some boundary equilibria (besides the origin). However, these non-zero boundary equilibria can be considered as a positive equilibrium of a subsystem, the one restricted to the boundary. Moreover, we can also use the solution of the polynomial complementarity problem to find some potential stable boundary equilibria. Thus, all the results and techniques in this section are also applicable to that case.
4 Cooperative Lotka-Volterra Model